OpenStudy (anonymous):

simplify the expression n^2-2n/n^2-3n

OpenStudy (anonymous):

Assuming your problem looks like \[\frac{n^2-2n}{n^2-3n}\] You can first factor an n out of both the top and bottom \[\frac{(n)n-2}{(n)n-3}\] And then cancel them, leaving you with a nice, neat \[\frac{n-2}{n-3}\] It's worth noting that the original equation is undefined at n=0, and both forms are undefined at n=3. It has a single root at x=2.

OpenStudy (anonymous):

so what would be the final form

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That would be the final form. There are a number of others you could get to, like \[\frac{1}{n-3}+1\]but there's little reason to do so.